OpenStudy (anonymous):

Determine what values of x the expression 5-x (and 5-x is inside a radical)

OpenStudy (anonymous):

i need some help please?

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

i don't understand how to solve the question, could you help please?

OpenStudy (anonymous):

What equation? You only provided an expression of \[\sqrt{5 - x}\]

OpenStudy (anonymous):

that is the equation, but the question is asking me to determine the values of x in the expession

OpenStudy (anonymous):

Do you mean the range of the expression? There's no equation to solve. An equation should have an equal sign

OpenStudy (anonymous):

I meant domain lol

OpenStudy (anonymous):

well it's an expression and i have to determine the values of x in the expression and express my answer in an inequality

OpenStudy (anonymous):

Oh so they probably mean the domain of the function

jimthompson5910 (jim_thompson5910):

sounds like you want the domain you cannot take the square root of a negative number, so 5-x cannot be negative this means \[\Large 5-x\ge 0\] solve for x to get \[\Large 5-x\ge 0\] \[\Large 5-x+x\ge 0+x\] \[\Large 5\ge x\] \[\Large x\le 5\]

jimthompson5910 (jim_thompson5910):

So only values that are either equal to 5 or less than 5 are allowed to be plugged into x

OpenStudy (anonymous):

oh, that makes sense how do i solve 3x-1over 5 =8

jimthompson5910 (jim_thompson5910):

Is it \[\Large \frac{3x-1}{5} = 8\] or \[\Large 3x-\frac{1}{5} = 8\] ??